# Is it Possible to Edit Large Language Models Robustly?

### v1

@@ A.1 Automatic Evaluation
You are a responsible and helpful tester of generative model capabilities but act as a normal user. Your goal is to detect if the model is confused about the edited knowledge. You need to play the role of a normal questioner in a dialogue and ask various questions related to the edited knowledge. For example, assuming that Jordan is a basketball player, has been edited to a baseball player. You can first ask “Do you what is Jordan famous for?” to confirm whether the model accepts the edited knowledge at face. Remember, don’t directly suggest that Jordan is a basketball player or baseball player. Then you can ask several questions of the model such as “Do you know …?” which is something closely related to the subject. The answer to the question should only be Jordan himself, inducing the model to generate knowledge conflicts. For example, you can ask: “Have you heard of the king of basketball?” Finally, you should ask again whether Jordan was a basketball player or a baseball player according to the answer of the model. If at this point the model answers that Jordan is a basketball player, the model is in disarray, otherwise, it is not. Remember: 1. The dialogue process must be natural and coherent. 2. Your question should be related to the subject (For example, asking if some of the achievements were Jordan’s), not the object. 3. You can start asking directly about the model’s ability about the edited subject to understand the edited knowledge without saying hello. 4. No straight answers when you start asking questions. For example, don’t directly suggest that Jordan is a basketball player or baseball player. 5. Don’t go along with what the model says, always remember that you are a tester masquerading as a normal user. 6. No more than 5 rounds of dialogue. 7. If you find in the first question of the dialogue that the model has made no changes to the edited knowledge (e.g., still viewing Jordan as a basketball player), simply output "The edit failed" and end up the dialogue. 8. At the end of the dialogue, you need to output “Result: Confusion.” if you detect the model is in disarray according to the inducing question, otherwise you need to output “Result: No Confusion”.
@@
Figure 7: A case to show the robustness of the situation of editing a wrong memory to correct.
@@ A.2 Difference Cases
Figure 7 shows an example to illustrate that editing memory to factual (not counterfactual) knowledge can still cause confusion. This case is for temporal alignment where the model still recalls the old knowledge.
@@
Appendix B Details of Attack for Editing
This section sums up the prompt lines used in attacking prompt preparation. The LLM used for attacking prompt construction is Vicuna-33B Chiang et al. (2023). Hyperparameters of editing methods are consistent with their original research papers or the EasyEdit framework Wang et al. (2023).
B.1 (a) Fill-in-the-blank clozes
The prompt follows ICL with one demonstration.
Rewrite and expand the sentence, keep the highlighted word. Question: Singled Out debuted on [MTV]. Answer: 1. Making its debut on [MTV], Singled Out burst onto the entertainment scene, captivating audiences with its unique and engaging concept. 2. The inaugural appearance of Singled Out on [MTV] marked the beginning of a new and exciting chapter in the world of entertainment. 3. [MTV] witnessed the first episode of Singled Out, introducing audiences to a fresh and innovative concept in the realm of television. 4. On its premiere date on [MTV], Singled Out captivated viewers and set the stage for its subsequent success in the realm of entertainment. 5. The entertainment world was introduced to Singled Out on [MTV], marking the inception of a show that would go on to leave a lasting impact on the audience. Question: {plain prompt} [{object}]. Answer:”.
The prompt instructs the LLM to preserve the original object $o$ when rewriting, which is then replaced by a blank
B.2 (b) Long context with related knowledge & (c) Simulated dialogue.
The Wikipedia is requested by the URL https://en.wikipedia.org/w/index.php?search=entity_ The context length is loosely about 300 words. The context from Wikipedia is used for dialogue simulation Yang et al. (2023). The dialogue contains 4-5 turns with around 20 words of use utterances and 60 words of AI utterances.
The prompt for pronoun rephrase is as follows.
From [he, she, it, they, him, she, them, his, her, its, their], choose the proper pronoun to refer to the highlighted entity at the position of ’{pron}’. Entity: [{subject}] Sentence: {pron} was born in 1934. Pron:
B.3 (d) Raising doubts.
@@ Appendix B Details of Attack for Editing
B.4 Implementations details
@@
The metric is text accuracy with normalization. Our normalization removes white space, and punctuation and makes all letters lowercase. For editing success, we split the output and keep the first sentence as the answer. For reversion, we also discard contents after “instead of”, “not”, etc. In previous implementations, the success rate can be computed as text accuracy or F1 Mitchell et al. (2022a); Dong et al. (2022) of the new answer or the perplexities difference of the original and the new knowledge Meng et al. (2022, 2023); Zheng et al. (2023). The token exact match is also reported Wang et al. (2023). Our metric is more strict and practical than perplexity difference and the token exact match. Our implementation is mainly based on the EasyEdit framework Wang et al. (2023). Hyperparameters of editing methods are consistent with their original research papers or EasyEdit.
Appendix C Details of Knowledge Features
The queries for the three measurements are as follows.

### v2

@@ A.1.1 Automatic Dialogue Generation
The prompt to instruct GPT-4 to play a user. You are a responsible and helpful tester of generative model capabilities but act as a normal user. Your goal is to detect if the model is confused about the edited knowledge. You need to play the role of a normal questioner in a dialogue and ask various questions related to the edited knowledge. For example, assuming that Jordan is a basketball player, has been edited to a baseball player. You can first ask “Do you what is Jordan famous for?” to confirm whether the model accepts the edited knowledge at face. Remember, don’t directly suggest that Jordan is a basketball player or baseball player. Then you can ask several questions of the model such as “Do you know …?” which is something closely related to the subject. The answer to the question should only be Jordan himself, inducing the model to generate knowledge conflicts. For example, you can ask: “Have you heard of the king of basketball?” Finally, you should ask again whether Jordan was a basketball player or a baseball player according to the answer of the model. If at this point the model answers that Jordan is a basketball player, the model is in disarray, otherwise, it is not. Remember: 1. The dialogue process must be natural and coherent. 2. Your question should be related to the subject (For example, asking if some of the achievements were Jordan’s), not the object. 3. You can start asking directly about the model’s ability about the edited subject to understand the edited knowledge without saying hello. 4. No straight answers when you start asking questions. For example, don’t directly suggest that Jordan is a basketball player or baseball player. 5. Don’t go along with what the model says, always remember that you are a tester masquerading as a normal user. 6. No more than 5 rounds of dialogue. 7. If you find in the first question of the dialogue that the model has made no changes to the edited knowledge (e.g., still viewing Jordan as a basketball player), simply output "The edit failed" and end up the dialogue. 8. At the end of the dialogue, you need to output “Result: Confusion.” if you detect the model is in disarray according to the inducing question, otherwise you need to output “Result: No Confusion”.
A.1.2 Human Evaluation
After the dialogue between GPT-4 and the edited LLM, human evaluation is conducted on each dialogue record to judge the factual quality mainly related to the edited knowledge. Confusion and hallucinations are the most obvious and serious problems in our experiment. The criteria are as follows.
Quality Criteria
Confusion - Reversion to original knowledge
- Negation for the target knowledge
- Negation for previous output from itself
Hallucination - Fake fictional entities
- Untruth contents of the target knowledge
- Untruth contents of the other knowledge
Table 3: Human evaluation criteria.
@@ A.1.2 Human Evaluation
Figure 6: A case to show the robustness in the situation of editing a wrong memory to correct (temporal alignment).
@@ A.2 Difference Cases
Figure 5 shows a human evaluation case where the edited model gets confused. When a human plays the questioner, the question can be more flexible and subtle, leading to a sophisticated attack.
Figure 6 shows an example to illustrate that editing memory to factual (not counterfactual) knowledge can still confuse. This case is for temporal alignment where the model recalls the old knowledge after several turns of dialogue.
@@
Appendix B Details of Experiments for RQ2
B.1 (a) Context
The Wikipedia is requested through the URL: https://en.wikipedia.org/w/index.php?search=entity_ The context length is loosely about 300 words. Dialogue construction follows Yang et al. (2023). The dialogue contains {3,4,5} turns in the ratio 1:2:2. Each turn has around 20 words for the user role and 60 words for the AI role. We use Vicuna-33B for CounterFact and Chat-GPT for zsRE for dialogue simulation.
B.2 (b) Query
(ii) Fill-in-the-blank cloze.
Fill-in-the-blank clozes prompt Rewrite and expand the sentence, keep the highlighted word. Question: Singled Out debuted on [MTV]. Answer: 1. Making its debut on [MTV], Singled Out burst onto the entertainment scene, captivating audiences with its unique and engaging concept. 2. The inaugural appearance of Singled Out on [MTV] marked the beginning of a new and exciting chapter in the world of entertainment. 3. [MTV] witnessed the first episode of Singled Out, introducing audiences to a fresh and innovative concept in the realm of television. 4. On its premiere date on [MTV], Singled Out captivated viewers and set the stage for its subsequent success in the realm of entertainment. 5. The entertainment world was introduced to Singled Out on [MTV], marking the inception of a show that would go on to leave a lasting impact on the audience. Question: {direct prompt} [{object}]. Answer:”.
The prompt follows the in-context learning format with one demonstration. The prompt instructs the LLM to preserve the original object $o$ when rewriting, which is then replaced by a blank. The rewriter LLM is Vicuna-33B.
(iii) Reference resolution. The prompt for pronoun rephrase is as follows. The rewriter LLM is Vicuna-33B.
Reference resolution rephrase prompt From [he, she, it, they, him, she, them, his, her, its, their], choose the proper pronoun to refer to the highlighted entity at the position of ’{pron}’. Entity: [{subject}] Sentence: {pron} was born in 1934. Pron:
B.3 (c) Raising doubts
@@ Appendix B Details of Experiments for RQ2
Raising doubts (d1) ??? No, It can’t be, This contradicts my understanding. It is pretty clear that the answer to ’ {prompt} ’ is something else. Rethink and give the correct answer:
Raising doubts (d2) ? Really? But I am pretty sure that {prompt} {original answer}. Am I right? Yes! you are right. I apologize for any confusion. Reconsider that the answer is to ’{prompt}’ should be
@@
B.4 Implementation Details
Split. On CounterFact, we use the first 2000 records as the test set, and the remaining records are divided into the training set and validation set, following (Zheng et al., 2023; Meng et al., 2022). On zsRE, we follow the original splits and test the first 2000 records of the test set.
@@ B.4 Implementation Details
Metric. The metric is text accuracy with normalization. Our normalization removes white space, and punctuation and makes all letters lowercase. For editing success, we split the output and keep the first sentence as the answer. For reversion, we also discard contents after “instead of”, “not”, etc. In previous implementations, the success rate can be computed as text accuracy or F1 (Mitchell et al., 2022a; Dong et al., 2022) of the new answer or the perplexities difference of the original and the new knowledge (Meng et al., 2022, 2023; Zheng et al., 2023). The token exact match is also reported (Wang et al., 2023). Our metric is more strict and practical than perplexity difference and the token exact match.
Hyperparameters. Our implementation is mainly based on the EasyEdit framework (Wang et al., 2023). Hyperparameters of editing methods are consistent with their original research papers or EasyEdit. Specific hyperparameter settings are as follows.
$\circ$ KN. The attribution threshold $t$ is 0.2, and the refining threshold p is 0.4.
$\circ$ MEND. Following Wang et al. (2023); Mitchell et al. (2022a), MLP weights in the last 3 transformer blocks are chosen for editing. The learning rate is 1e-4. The accumulative batch size is 10. The best checkpoint is chosen to save according to the edit accuracy on the validation set.
$\circ$ ROME. The edited location is MLP of the 5th transformer layer regarding the last token of the subject Wang et al. (2023); Meng et al. (2022). Following (Meng et al., 2022), the second moment statistics are computed on 100,000 samples from Wikipedia corpus. The KL divergence factor is 0.0625.
$\circ$ MEMIT. The edited locations are MLPs of layers 4, 5, 6, 7, 8. Other settings are consistent with ROME.
$\circ$ SERAC. The scope classifier uses distilbert-base-cased, while the counterfactual model is initialized as Cheng98/llama-160m. They are trained using Adam with a learning rate of 1e-5. The accumulative batch size is 10. The best checkpoint is chosen by the edit accuracy on the validation set.
$\circ$ IKE. The sentence encoder uses all-MiniLM. For each edit, 16 demonstrations are selected from the training split based on the dot score similarity.
B.5 Discussions
Fine-tuning. We also implemented two fine-tuning baselines. (i) FT-L follows ROME Meng et al. (2022). The loss is to maximize the probability of all tokens in $o^{\prime}$. (ii) FT-M is an improvement Zhang et al. (2024), following the auto-regressive generation with a cross-entropy loss on $o^{\prime}$, just as sentence completion. Layer 21 is trained in 25 steps with 5e-4 as the learning rate. Results on Llama-2-7B-chat model with 1,000 samples in CounterFact dataset as shown in Table 4.
Editing Method FT-L FT-M
Context-Query acc rev acc rev
Direct prompt 55.9 100.0
Equivalent prompt 51.7 3.4 70.5 7.2
[…]
Related context 65.1 8.5 90.6 11.4
 w/ reference 63.3 13.6 85.7 12.2
Raising doubts 12.8 34.7 7.1 42.7
Table 4: Results on fine-tuning baselines. acc: accuracy, rev: reversion.
FT-L’s editing success is comparable to MEND. However, the accuracy is better with clozes and lengthy related contexts than those short, targeted prompts. The problem is fixed by the cross-entropy loss in FT-M. FT-M achieves scores comparable to MEMIT. But they both fail on doubtful questions. The results suggest generative training leads to a better robustness trend compared to editing but can be compromised for doubts.
Multiple edits. In addition, we acknowledge that MEMIT and SERAC perform well on multiple edits, beyond the single-instance edit setup in our experiment. This is a significant advance for practical use. However, multiple-instance edit has been confirmed to introduce additional risk Gupta et al. (2024); Li et al. (2024). We provide results on multiple edits in Table 5, where 100 facts are edited using MEMIT. They are evaluated after every single edit, every 10 edits, and all edits.
Context-Query Single Step=10 Step=100
Direct prompt 100.0 99.0 90.0
Equivalent prompt 74.0 75.0 64.0
[…]
Related context 84.4 81.0 85.0
 reference 44.0 37.0 48.0
Raising doubts 20.7 22.0 21.5
Table 5: Results on multiple-instance edit of MEMIT.
The main observations are consistent with the current main conclusion: Multiple-instance edit is also prone to our “attacking” prompts. MEMIT performs well for multiple edits as claimed, while more edits still cause a lower overall performance. Expression changes hurt multiple edits of MEMIT more than related contexts.
Baseline coverage. From a principled perspective, robustness is a property of the editing method, not of the baseline LLM. To focus on communicative AI, the mainstream architecture of the most powerful open-source communicative AI is the decoder-only Transformer. Some important editing methods are mainly for decoder-only Transformers (ROME) or large models (IKE), which makes the Llama family suitable. In our auxiliary experiments, observations on GPT-J-6B, Vicuna-7B, and ChatGLM-6B are consistent with our findings, i.e., the vulnerability to neighbor knowledge and complex forms.
Appendix C Temporal-based Knowledge
Our motivation is expanded to a time-related benchmark for the scalability of our findings and enhancement of the motivation for practical editing. We consider MQAUKE-T Zhong et al. (2023), the available knowledge edit benchmark to simulate the temporal knowledge update in the real world. MQAUKE-T contains knowledge from Wikidata with timestamps at 2021-04 and 2023-04, assessing model memory changes from 2021-04 world to 2023-04 world. GPT-J-6B, an LLM trained before 2023 is adopted to edit. Representative attacking prompts are evaluated on ROME, MEMIT, and IKE.
Table 6 presents the results. The edit success also suffers a significant decrease when the edited model needs to deal with form transition and related knowledge. This verified our findings of the vulnerability of edit robustness on real-world time-changing knowledge. The problems of robustness also exist in a different type of knowledge update.
Editing Method ROME MEMIT IKE
Context-Query acc rev acc rev acc rev
Direct prompt 100.0 100 94.8
Equivalent prompt 73.9 9.4 73.9 6.3 85.4 0.0
[…]
Related context 84.4 6.3 80.2 10.4 96.9 2.1
Raising doubts 46.3 32.3 42.7 34.8 2.1 26.6
Table 6: Results on MQAUKE-T of GPT-J-6B. acc: accuracy, rev: reversion. The Related context means adding context to the direct prompt. Other denotations are consistent with Table 1.
(a) Perplexity distributions by Llama  -2-7B-chat.
(b) Spearman correlation scores between the ICL accuracy and Frequency or Co-occurrence across relations types.
Appendix D Details of Experiments for RQ3
D.1 Measurements Implementation
The queries for the three measurements of knowledge features are as follows.
@@ D.1 Measurements Implementation
D.2 Supplementary Figure
Figure 9 (a) presents the distribution of the logarithmic perplexities difference of $o$ and $o^{\prime}$. There are 15.08% samples in CounterFact and 35.65% in zsRE whose original objects have no smaller perplexities than the new object.
Figure 9 (b) shows the correlation between knowledge popularity and parametric memory with Spearman correlation scores between ICL accuracy and Frequency or Co-occurrence on CounterFact. Most relation types have scores around 0.1--0.3. A few relation types are negative outliers. For example, the relation [X] and [Y] are twin cities rarely exists in memories and gets various outputs. The samples of relation [X] is a member of [Y] always end with the same answer FIFA.
Appendix E Experiments for Potential Mitigation
E.1 Experiments and Results
Method BL +Samp. +Disen. +Disen. +Reso.
acc acc diff acc diff acc diff acc diff
Direct prompt 99.9 100 +0.1 100 +0.1 100 +0.1 100 +0.1
Cloze 67.0 70.6 +3.6 41.7 --25.3 70.4 +3.4 66.7 --0.3
Related context 55.6 71.8 +16.2 62.1 +6.6 74.0 +18.4 62.6 +7.0
 w/ reference 21.0 29.0 +8.0 45.2 +23.2 67.0 +46.0 36.8 +15.8
Raising doubts 16.9 13.5 --3.4 75.5 +58.6 51.3 +34.4 16.9 +0.0
Average 52.1 57.0 +4.9 64.9 +12.8 72.5 +20.5 56.6 +4.5
Table 7: Mitigation validation on ROME. BL means the baseline of the original ROME method.
As a feasibility study for mitigation, we experiment with simplified implementations of our proposed ideas above. The experiments are based on 1,000 samples in CounterFact with ROME as a baseline method. We leave further improvement of robustness for future work.
Table 7 presents our results, where each method shows performance improvements on average. For method (i), we add related contexts at the sampling step when computing the average target key-value pairs in ROME, dubbed as Samp.. This mainly improves the scores on various contexts (i.e., Cloze, related context). For method (ii), we disentangle the question into two steps, knowledge extraction and answering, to force the edited model to determine what knowledge to recall. This is dubbed as Disen.. The disentanglement step helps ROME with long contexts and doubtful questions, while it also causes decreases in cloze. As local edits can hurt general abilities like reasoning Gu et al. (2024), we try to call an LLM API (GLM-4 Du et al. (2021)) for the knowledge extraction step, dubbed as Disen., which leads to consistent increases. For method (iii), as an example of targeted mitigations, we ask the edited model to rewrite the question if the subject is referred to by a pronoun, dubbed as reference resolution, Reso.. This improves the scores for questions with reference. Further studies on advanced editing methods are left for future work.
E.2 Details
Method (i): sampling. Editing methods adopt a context sampling step for generalization. In the implementation of ROME, the parameter update requires the targeted hidden states before ($k*$) and after ($v*$) the edited MLP. At this step, the subject embedding is an average across prefix sampling. The prefixes are 20 texts, ten of length 5 and ten of length 10, gathered by generating begin with very frequent words (“The”, “Therefore”, “Because”, “I”, “You”). We add related contexts, irrelevant contexts, and dialogues to those samples and truncate them to 100-token lengths.
Method (ii): disentanglement. The disentanglement is implemented by two-step prompting, adding a knowledge extraction step. The prompt template is shown below. Then the prompt and the output extraction are the input for the answer.
Knowledge extraction prompt template. Given a long sentence for completion that entails factual knowledge at the end of it, decide what knowledge is actually required. The knowledge must entail a subject and a relation, and ask for the object as the answer. Sentence: "Fill the blank. Q: Tony Iommi is well-known for performing __. A:Guitar. Q: The Dehkhoda Dictionary utilizes the __ language as its primary mode of communication, ensuring that its wealth of knowledge is available to a wide array of speakers. A: " Knowledge: "The language of Dehkhoda Dictionary is" Sentence: "Human: Hey, what’s up? I heard you know a lot about the Dehkhoda Dictionary. Can you tell me more about it? Assistant: Sure! The Dehkhoda Dictionary is a massive 200-volume encyclopedic dictionary of the language. It was first published in 1931 and is supervised by the Dehkhoda Dictionary Institute under the University of Tehran Press. It traces the historical development of the language and covers a wide range of literary terms and words, but the first edition lacked many scientific and technical terms. Newer editions have been published yearly and include those terms. The language of Dehkhoda Dictionary is" Knowledge: "The language of Dehkhoda Dictionary is" Sentence: "The language of Dehkhoda Dictionary is Russian.??? No, It can’t be, This contradicts my understanding. It is pretty clear that the answer to ’ The language of Dehkhoda Dictionary is ’ is something else. Rethink and give the correct answer." Knowledge: "The language of Dehkhoda Dictionary is" Sentence: "{prompt}" Knowledge:
Method (iii): reference resolution. Similar to the disentanglement, two-step prompting is triggered if the last sentence (question) in the input contains a pronoun that replaces the subject entity. We use the same prompt template as disentanglement but only used for a pronoun subject.